Bond J has a coupon rate of 4 percent. Bond K has a coupon rate of 14 percent. Both bonds have 17 years to maturity, a par value of $1,000, and a YTM of 8 percent, and both make semiannual payments. 

a. If interest rates suddenly rise by 2 percent, what is the percentage change in the price of these bonds? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.) 

b.If interest rates suddenly fall by 2 percent instead, what is the percentage change in the price of these bonds? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.) 
Answer #1
a. Bond J -18.61%
Bond K -14.72%
Working:
# 1 Existing price of Bond J = =-pv(rate,nper,pmt,fv) Where,
= $     631.78 rate = 8%*(6/12) = 0.04
nper = 17*(12/6) = 34
pmt = 1000*4%*(6/12) = $       20.00
fv = $ 1,000.00
Existing price of Bond K = =-pv(rate,nper,pmt,fv) Where,
= $ 1,552.34 rate = 8%*(6/12) = 0.04
nper = 17*(12/6) = 34
pmt = 1000*14%*(6/12) = $       70.00
fv = $ 1,000.00
# 2 Price after rise of interest rate by 2 percent:
New price of Bond J = =-pv(rate,nper,pmt,fv) Where,
= $     514.21 rate = 10%*(6/12) = 0.05
nper = 17*(12/6) = 34
pmt = 1000*4%*(6/12) = $       20.00
fv = $ 1,000.00
New price of Bond K = =-pv(rate,nper,pmt,fv) Where,
= $ 1,323.86 rate = 10%*(6/12) = 0.05
nper = 17*(12/6) = 34
pmt = 1000*14%*(6/12) = $       70.00
fv = $ 1,000.00
# 3 Percentage change in the price of:
Bond J = (b-a)/a Where,
= -18.61% a Existing price = $     631.78
b New Price = $     514.21
Bond K = (b-a)/a Where,
= -14.72% a Existing price = $ 1,552.34
b New Price = $ 1,323.86
b. Bond J 24.84%
Bond K 18.87%
Working:
# 1 Price after fall of interest rate by 2 percent:
New price of Bond J = =-pv(rate,nper,pmt,fv) Where,
= $     788.68 rate = 6%*(6/12) = 0.03
nper = 17*(12/6) = 34
pmt = 1000*4%*(6/12) = $       20.00
fv = $ 1,000.00
New price of Bond K = =-pv(rate,nper,pmt,fv) Where,
= $ 1,845.27 rate = 6%*(6/12) = 0.03
nper = 17*(12/6) = 34
pmt = 1000*14%*(6/12) = $       70.00
fv = $ 1,000.00
# 2 Percentage change in the price of:
Bond J = (b-a)/a Where,
= 24.84% a Existing price = $     631.78
b New Price = $     788.68
Bond K = (b-a)/a Where,
= 18.87% a Existing price = $ 1,552.34
b New Price = $ 1,845.27